OpenStudy (anonymous):

Use the compound interest formula to compute the total amount accumulated. $10,000 for 20 years at 5% compounded annually A. $20,000.00 B. $25,269.50 C. $19,500.00 D. $26,532.98

OpenStudy (anonymous):

An = Ao ( 1 + r )^y or An = 10000 ( 1 + .05 )^(20) = $ 26,532.98